Given that the regression equations of Y on X and of X on Y are respectively Y = X
and 4X − Y = 3, and the second moment of X about the origin is 2, find
i) the means of X and Y;
ii) the correlation coefficient between X and Y;
iii) the standard deviation of Y
1
Expert's answer
2019-03-25T15:32:41-0400

(i) Since two regression lines always intersect at a point representing mean values


"(\bar{x}, \bar{y})""Y=X""4X-Y=3"

Then


"\bar{x}=1, \bar{y}=1"

(ii) To find the given regression equations in such a way that the coefficient of dependent variable is less than one at least in one equation. So


"4X-Y=3\implies4X=3+Y \implies X={3 \over 4}+{1 \over 4}Y"

That is


"b_{xy}={1 \over 4}=0.25"

"Y=X"

That is


"b_{yx}=1"

Hence coefficient of correlation r between x and y is given by:


"r=\sqrt{b_{xy}*b_{yx}}=\sqrt{0.25*1}=0.5"

(iii) To determine the standard deviation of y , consider the formula:


"\sigma_y={{b_{yx}*\sigma_x} \over r}"

"{\sigma_x}^2=E[X^2]-(E[X])^2=2-(1)^2=1"

"\sigma_x=1"

"\sigma_y={1*1 \over 0.5}=2"